OpenStudy (anonymous):

A rocket is launched from atop a 54-foot cliff with an initial velocity of 119 feet per second. The height of the rocket t seconds after launch is given by the equation h = –16t 2 + 119t + 54. Graph the equation to find out how long after the rocket is launched it will hit the ground. Estimate your answer to the nearest hundredth of a second.

OpenStudy (campbell_st):

well to hit the ground, the height = 0 so you have \[0 = -16t^2 + 119t + 54\] use the general quadratic formula to find t. and only use the positive t value in your answer

OpenStudy (anonymous):

i dont know the quadratic formula well

OpenStudy (campbell_st):

ok... its \[t = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] you have a = -16, b = 119 and c = 54 substitute the values and calculate the values of t. only use the positive value hope it helps